OpenStudy (anonymous):

A student is attempting to find the equation of a line given a point and the slope. Which of the following statements best applies to the sample mathematical work shown below? Given (10, –2) with a slope of –1, I find an equation for the line by plugging in values I'm given into a point-slope form equation. Thus I have y – (–2) = –1( x –10), which results in an equation of y = –x + 8.

OpenStudy (anonymous):

Given (10, –2) with a slope of –1. y + k = a(x + h) (-2) + k = -((10) + h) k = -8 - h ... cannot be solved

OpenStudy (anonymous):

y = ax + b (-2) = -(10) + b b = 8

OpenStudy (texaschic101):

y - (-2) = -1(x - 10) y + 2 = -x + 10 y = -x + 10 - 2 y = -x + 8 um...is there answer choices....because this statement is correct

OpenStudy (anonymous):

so then you get y = -x + 8

OpenStudy (anonymous):

The student improperly set up the equation to find the y-intercept. B. The student used the correct equation to find the y-intercept. However, he did not solve for the y-intercept properly. C. The slope in the final equation is not correct. D. The mathematical work shown above is correct.

OpenStudy (texaschic101):

correct...although it could have been solved with just the slope intercept formula

OpenStudy (anonymous):

soo its D?